OpenStudy (anonymous):

Find the derivative of 2x^2-13x+5 and use it to find an equation of the line tangent to the curve at x=3

OpenStudy (anonymous):

Well, first you need to substitute x into f(x) to get a y value. Then derive the f(x) to get the equation for the gradient) (I'm trying to help you rather than do it for you lol)

OpenStudy (anonymous):

okay I got -16 so its (3,-16) and then I take the limit of f(x)? thank you by the way

OpenStudy (anonymous):

Uh.. I'm not sure how you got -16. Substitute x=3 into 2x^2 - 13x + 5 So it will be (2 * (3^2)) - (13 * 3) + 5...

OpenStudy (anonymous):

what are you getting? I still got -16

OpenStudy (anonymous):

(2 * 9) - 29 + 5 18 - 29 + 5 -11 + 5 -6

OpenStudy (anonymous):

13 x 3 is 39

OpenStudy (anonymous):

OH. Jeez. That's embarrassing. *cough* let's just forget that happened, eh? Alright, -16 is your y value. So the coordinates you want to find the equation to the tangent of are (3,-16). Now derive your f(x) = 2x^2 - 13x + 5

OpenStudy (anonymous):

lol its cool.. I got 2x-13 is that correct?

OpenStudy (anonymous):

Almost, not quite.. When you derive a function... f(x) = x^n f'(x) = nx^(n-1) You multiply the coefficient by the power and subtract one from the power. You missed something ;)

OpenStudy (anonymous):

Oh yeah okay so now we need to find what f'(3)?

OpenStudy (anonymous):

Yeah. f'(x) is your gradient function, ie it will give you the gradient of the main function at any point. You want the gradient at (3,-16), so sub in 3 for x and that will give you the gradient at that point. So what's your fixed derived equation? :)

OpenStudy (anonymous):

f'(x)=4x-13 f'(3)=4(3)-13 f'(3)=-1 y+16=-1(x-3) then I get....

OpenStudy (anonymous):

Exactly! Now just expand and rearrange for y.

OpenStudy (anonymous):

y=-1x-13

OpenStudy (anonymous):

Yup! That's the equation for the tangent at that point. You can double check with a graphics calculator or similar to make sure the lines meet at that point.

OpenStudy (anonymous):

looks right :) Thank you so much

OpenStudy (anonymous):

No worries dude :) have a great day

OpenStudy (anonymous):

you too

OpenStudy (anonymous):

Legend <3

OpenStudy (anonymous):

Can I ask one more question?

OpenStudy (anonymous):

Sure thing man, go ahead :)